OpenStudy (love_to_love_you):

Which equations represent hyperbolas? Please help, will fan and medal!

OpenStudy (love_to_love_you):

2x^2 + 16x + 2y^2 + 14y − 9 = 0 2x^2 + 4x − 5y^2 − 10y + 57 = 0 -x^2 + 5x − 7y^2 + 2y − 81 = 0 x − 2y^2 + 4y + 15 = 0 -x^2 + 12x + 3y^2 + 7y + 11 = 0

OpenStudy (love_to_love_you):

@jim_thompson5910

OpenStudy (bobo-i-bo):

Okay... this is a bit complicated to fully explain. Hmm, firstly, do you know how to complete the square?

OpenStudy (agent0smith):

Any equation which HAS both an x^2 and a y^2 term, but both are NOT the same sign.

OpenStudy (love_to_love_you):

I know nothing about this

OpenStudy (agent0smith):

Everything you need is in my last post.

OpenStudy (love_to_love_you):

Feeling really stupid, but I still dont get it

OpenStudy (bobo-i-bo):

The only stupid thing is to think that you're stupid. :P

OpenStudy (bobo-i-bo):

Take a look at this: https://www.mathsisfun.com/algebra/completing-square.html

OpenStudy (bobo-i-bo):

It's very important thing to learn tbh!

OpenStudy (bobo-i-bo):

Okay, so let me give you an example

OpenStudy (love_to_love_you):

But how does that show me which ones represent hyperbolas? I'm completely new to this course so I'm extremely confused

OpenStudy (bobo-i-bo):



OpenStudy (bobo-i-bo):



OpenStudy (bobo-i-bo):

That's the general shape of a hyperbola

OpenStudy (agent0smith):

Any equation which HAS both an x^2 and a y^2 term, but both are NOT the same sign. Eg. 2x^2 + y - 5y^2 - 6x = 0 IS a hyperbola because it has BOTH an x^2 and a y^2 term, but one is positive and one is negative, so they aren't the same sign.

OpenStudy (bobo-i-bo):

Give me a sec :P

OpenStudy (agent0smith):

@Bobo-i-bo you don't need to complete the square to tell if they're hyperbolas or not.

OpenStudy (love_to_love_you):

Ohhh, okay.

OpenStudy (bobo-i-bo):

No indeed you don't, but if you want to fully understand hyperbolas and stuff, it's better... it's better to give a man a net rather than one fish afterall. :P

OpenStudy (love_to_love_you):

So the hyperbolas would be the second and last equations

OpenStudy (bobo-i-bo):

Play around with this: https://www.desmos.com/calculator/jvt3ahfmix

OpenStudy (agent0smith):

Yes @Bobo-i-bo , in general you need to know how to do other things to deal with them, but not to identify whether it's a hyperbola - in that case it's better to just know how to identify them.

OpenStudy (agent0smith):

@love_to_love_you correct. The second and last are the only hyperbolas.

OpenStudy (bobo-i-bo):

Currently it is a circle. Now try setting either a or b to be negative. Play around with the values :)

OpenStudy (bobo-i-bo):

@agent0smith I guess :P But personally understanding is more important to me.

OpenStudy (agent0smith):

@Bobo-i-bo understanding what? Understanding how to identify them is what's important here. That's a whole section of conics - knowing how to identify just by looking at the equation. @love_to_love_you if you really have no idea about this topic, I'd find it on khanacademy... you aren't going to learn the entire topic here on openstudy. There's a lot to learn about hyperbolas, as well as ellipses and parabolas.

OpenStudy (agent0smith):

Completing the square is not the way to identify a conic. That's far too tedious to identify a conic. Ideally, a student should understand that long before they get to the point of identifying a conic section from an equation (which is meant to be after they've learned about all the conic sections). "personally understanding is more important to me"... makes it seem like you didn't realize there was an alternative method to identify conics.

OpenStudy (love_to_love_you):

It's alright, I think we're going to get a tutor for this, in the meantime I just need help on a few questions

OpenStudy (agent0smith):

I'd recommend that, this is a pretty tough area of algebra 2, and it's easier to learn about these in person when someone can show it to you easily.